If you invest $1,907 today at an interest rate of 8.51 percent, compounded daily, how much money will you have in your savings account in 27 years?
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

n = 27 * 365 days = 9855 days

r = 8.51%/365 (per day) = 0.0233% (per day)

FV = $1907 * (1 + 0.000233)9855

FV = $1907 * 9.9486

FV = $18,971.99